I've had the output from a fiber projected about 400mm onto the Hartmann sensor for around 5 days now. (The divergence angle from the fiber is around 86 mrad).
I played around with the temperature of the lab to induce some defocus changes in the Hartmann sensor. The system is mostly linear, but there are relatively frequent jumps in the defocus of the order of 1E-4 m^-1. This may be due to a number of things - the Hartmann plate may be moving, the fiber holder may be shifting back and forth, there may be some issue with the source wavelength shifting.
Sun 30th May 2011 - 11:40AM - the z-axis control on the NewFocus 9091 fiber coupling mount was not tightened. I tightened that to secure the control.